larry mitchell invested part of his $34,000 advance at 5% annual simple interest and the rest at 7% annual simple interest. the yearly total from both was $2,300. Find the amount at each rate.
Answer by fractalier(6550)

Call the amount invested at 5%, x.
Then the amount invested at 7% would be 34000-x.
The set up then looks like this...
.05x + .07(34000-x) = 2300
Now solve for x...
.05x + 2380 - .07x = 2300
-.02x = - 80
x = $4000 invested at 5% which means
34000-4000 = $30,000 invested at 7%